Calculate the molarity of the two solutions. The first solution contains 0.550 mol of NaOH in 2.95 L of solution. molarity: TCalculate the molality of a solution formed by adding 7.00 g NH,Cl to 14.0 g of water. molality: m NH C1What is the mole fraction of solute in a 3.69 m aqueous solution? Xsolute = 0.06An aqueous sodium acetate, NaC,H,O, , solution is made by dissolving 0.465 mol NaC,H,O, in 0.690 kg of water. Calculate the mAn aqueous potassium chloride, KCl, solution is made by dissolving 0.380 mol KCl in 500.0 g of water. Calculate the molality
Answer #1

Molarity of a solution is found using the equation: Molariy = Number of moles/ Volume in L

1. moles= 0.550 mol and V= 2.95L

Molarity = 0.550 moles/ 2.95L = 0.186 M

2. Given, ,mass= 19.1g and V= 627mL= 0.627 L

Number of moles= mass/ molar mass = 19.1g/ 58.44g/mol = 0.327 mol

Molarity = 0.327 mol/ 0.627L = 0.523 M

3. Molality = Number of moles/ Kg of solvent

Number of moles = 7g / 53.49 g/mol = 0.130 mol

Molality = 0.130 moles / 0.014kg = 9.28 m

4. 3.69 m solution means 3.69 moles of solute in 1 Kg of solvent

moles of solute =3.69

moles of solvent = 1000g/ 18g/mol = 55.55 mol

Mole fraction = moles of solute / moles of solute + moles of solvent= 3.69/ 3.69+55.55 = 0.062

5. molality = 0.465 mol / 0.690 kg = 0.673 m

6. molality = 0.380 mol/0.5Kg = 0.76 m